Ahmer walks 785m south and 95m west 760m north find the displacement and distance?
1
Expert's answer
2020-03-10T11:16:52-0400

Distance:


"d=785+95+760=1640\ m"

Displacement:


"s=\sqrt{(785-760)^2+95^2}=98\ m"

Direction:


"\theta=\arctan{\frac{785-760}{95}}=14.7\degree\text{S of W}"